OpenStudy (anonymous):

What are the solutions of the equation? z^2 - 6z + 5 = 0

OpenStudy (anonymous):

1 and 5

OpenStudy (anonymous):

(z-5)(z-1)=0

OpenStudy (anonymous):

z= 0 z = 11?

OpenStudy (anonymous):

z=1 or z=5

OpenStudy (anonymous):

factor it so (z-5)(z-1)=0 z-5=0 z-1=0

OpenStudy (anonymous):

ok i understand thanks

OpenStudy (anonymous):

Z=1 Z=5